OpenStudy (christos):

China has a population of 1350044605 people and a growth rate of 0.0439% per year. What's the 0.0439% of 1350044605 and how do I find it? Just the 0.0439%

OpenStudy (mertsj):

Do you know that in mathematics "of" means multiply?

OpenStudy (mertsj):

So change 0.0439% to a decimal and multiply it by 1,350,044,605

OpenStudy (christos):

Ok but, how would I go about finding the %? I modified the question so the of may not be the case here @Mertsj

OpenStudy (mertsj):

.0439%=.000436(1,350,044,605)=592669.5816

OpenStudy (anonymous):

Like this: (1350044605) * (0.0439/100) = (1350044605 * 0.0439) / 100 = 592669.5816